# Noise analysis of a quasi-phase-matched quantum frequency converter and higher-order counter-propagating SPDC

### v1

@@ Abstract
Quantum frequency conversion (QFC) will be an indispensable ingredient in future quantum technologies. For example, large-scale fibre-based quantum networks will require QFC to interconnect heterogeneous building blocks like emitters, channels, memories and detectors. The performance of existing QFC devices - typically realised in periodically-poled nonlinear crystals - is often severely limited by parasitic noise that arises when the pump wavelength lies between the wavelengths which are inter-converted. Here we investigate the noise spectrum of a frequency converter pumped by a CW 1064 nm laser. The converter was realised as a monolithic bulk pump enhancement cavity made from a periodically-poled potassium titanyl phosphate (ppKTP) crystal - quasi-phase-matched for the conversion of 637 nm to 1587 nm. In the range from 1140 nm to 1330 nm (up to 60 THz from the pump) Stokes-Raman resonances can be identified as the dominant noise source while the noise in the range from 1330 nm to 1650 nm can be attributed mainly to parasitic spontaneous parametric down-conversion (SPDC). Further, a succession of narrow-band peaks is observed in the spectrum originating from higher-order counter-propagating SPDC. Both types of counter-propagation, where either the lower-energy idler photon or the higher-energy signal photon counter-propagate relative to the pump beam, are observed, with narrow-band peaks corresponding to 10th to 44th order quasi-phase-matching.
@@ 1 Introduction
The efficient and noiseless frequency conversion of quantum light will be an essential building block in the development of future large-scale quantum networks. By enabling the coherent transfer of quantum information between disparate energies, it interfaces the low-loss capabilities of telecommunications infrastructure with the typically visible and near-IR photon energies required by matter-based qubits. One such example are nitrogen-vacancy (NV) centres in diamond [1, 2, 3]. Nitrogen-vacancy (NV) centers in diamond represent a promising platform due to the long coherence time of the electronic spin associated with this defect center [4], the possibility to couple this electronic spin to nuclear spin [5, 6, 7] and an optical interface at 637 nm [8]. With NV centers in diamond quite complex quantum networking tasks have been recently demonstrated [9, 10, 11]. Without the conversion of the 637 nm photons emitted by the NV center to telecommunication wavelengths, these photons would suffer dramatic transmission losses in optical fibres [12] that limit the size of these networks well below metropolitan scale. A variety of platforms are similarly limited by this, including other color centers in diamond like GeV (602 nm), SnV (620 nm), PbV (520/555 nm) [2] or quantum memory at 606 nm [13] and 580 nm [14], trapped ions (369.5 nm) [15] or hexagonal boron nitride (~600 nm) [16].
@@
In its simplest formulation, quantum frequency conversion (QFC) is realized as a difference- (sum-) frequency generation process in which a -nonlinear crystal converts a typically short wavelength photon to the telecommunications band (or vice versa) via a comparatively strong pump field at an intermediate wavelength. In this configuration, where the pump wavelength lies between the emitter and target wavelengths, the strong field required to ensure near-unity conversion efficiency also produces considerable unwanted noise at the target wavelength. This parasitic noise is not only present in the conversion from visible to telecommunication wavelengths, but also for the reverse process, since, for an efficient converter, any noise produced will itself be efficiently up- or down-converted. This parasitic noise – whether parametric or non-parametric in origin – has proven to be the limiting factor in the performance of these conversion devices.
@@
Efficient conversion demands a strength of non-linear interaction that is typically inaccessible in a three-wave-mixing process without either waveguide confinement [12], pump enhancement [17] or pulsed pump lasers [18]. Furthermore, frequency converters near-exclusively utilise quasi-phase matching, as it simultaneously allows for collinear interactions of the fields without spatial walk-off and unlocks access to the largest tensor elements of the nonlinear susceptibility. In theory, this allows for a dramatic reduction in pump power that ideally mitigates noise from non-parametric scattering of the pump photons, notably, Raman scattering and fluorescence. However, errors that arise in the periodic poling can vastly enhance the otherwise strongly-suppressed phase-mismatched parasitic spontaneous parametric down-conversion (SPDC) processes.
Furthermore, employing waveguide confinement to reduce the needed pump power can lead to a Cerenkov-idler configuration of the SPDC noise process [19] and waveguide inhomogeneities to further undesired phase-matching, both increasing the noise level even more.
@@
This emergent noise floor in the conversion between visible and telecommunication wavelengths was first reported by Albota and Wong [17] who considered conversion in a bulk periodically-poled lithium niobate (ppLN) crystal. Almost two decades on, converters based on ppLN waveguides, which form the current state-of-the-art, are still largely limited by pump-induced noise [20, 21, 22, 23, 24, 12]. The noise spectra of quantum frequency converters based such devices has been previously investigated in references [25, 24]. In the short-wavelength-pumped scheme of Ref. [24], parasitic SPDC is found to be the dominant source of noise. In the long-wavelength-pumped converter investigated in Ref. [25], anti-Stokes Raman scattering is identified as the main source of noise.
@@
A workaround for this noise problem is two-step conversion. Here the pump wavelength is chosen to be longer than the target wavelength (e.g. 2.1 ) and, due to energy conservation, no unwanted SPDC can be generated at the target wavelength. Additionally, the Raman shift for the anti-Stokes photon is very large, ensuring its noise contribution is negligible. First proposed in [17], two-step conversion was first demonstrated in a single waveguide device [26] and later in two separate waveguides [27]. A drawback of this approach is the necessity for a strong laser source in the mid-IR alongside the increased complexity of a two step process. Another approach to circumvent the parasitic noise arising in conversion is to utilise birefringent phase-matching in a bulk crystal [28], eliminating parasitic SPDC noise which would be enhanced by poling errors. Owing to the comparatively smaller nonlinear susceptibility and the effects of double refraction, this approach demands substantially more pump power and the maximum conversion efficiency is fundamentally bound below unity because of walk-off effects from double refraction.
@@
In our previous work we demonstrated an alternative route for low-noise QFC based on a single-step quantum frequency converter based on a monolithic cavity made from bulk periodically-poled potassium titanyl phosphate (ppKTP) [29]. The converter was designed for the conversion of single photons at a wavelength of 637 nm to the target wavelength in the telecommunication L-band (1587 nm) with a strong 1064 nm CW pump laser. Due to the improved poling quality of ppKTP [30, 31], parasitic noise arising from spurious SPDC was reduced by a factor of five compared to the best ppLN waveguide [12], while attaining a high internal conversion efficiency of 72% at about 72 W circulating pump power. The converter, however, was still exhibiting a pump-induced noise floor of 110 kHz/nm at the target wavelength. If quasi-phase-matched nonlinear crystals are to be employed for single-step converters between the visible and telecommunication wavelengths in realistic quantum network implementations considerable improvements in the noise floor are needed.
@@
In this work, we investigate the noise spectrum of the aforementioned bulk ppKTP quantum frequency converter designed to convert photons between 637 nm and 1587 nm via a strong 1064 nm pump laser. Our measurements fully characterise the resulting noise spectrum in the wavelength range of 1140 nm to 1650 nm. By resolving the polarisation and temperature dependence of features within the noise spectrum, we identify different noise contributions and the wavelength regions where they dominate. These results are of immediate relevance for most schemes where the strong pump field lies between the two interconnected wavelengths.
@@
Figure 1 provides a sketch of our conversion scheme and the primary noise contributions from the different processes. We find that close to this pump wavelength, Raman scattering of pump photons is the dominant noise source. As we consider increasingly longer wavelengths approaching the target wavelength, the required phonon energies increase and the contribution of Raman scattering to the noise background is quickly suppressed and spurious spontaneous parametric down-conversion (SPDC) becomes the dominant noise process. This is near-exclusively the case for quasi-phase-matched crystals, where errors in the periodic poling [32] drive spurious SPDC processes that produce a plateau in the noise-floor, effectively stable across the transparency of the material. This effect has been studied in detail in [20, 33, 34].
@@
In addition to the anticipated features from Raman scattering and the parasitic SPDC background, we have also identified an exotic nonlinear process - higher-order counter-propagating SPDC - that to our knowledge has never before been experimentally observed. This process, which is only present in periodically-poled devices, can contribute significantly to the parasitic noise background. Notably, both types of counter-propagation, where either the idler or signal photon counter-propagate relative to the pump beam, are observed, with narrow-band peaks of corresponding quasi-phase-matching orders m from 10 up to 44 visible. Owing to poling errors related to the duty-cycle, peaks associated with even orders of m are also visible.
@@
In totality, these results provide a comprehensive overview of the noise spectrum of a quantum frequency converter that employs a quasi-phase-matched process and a pump that lies in between the interconnected wavelengths. Our presented measurements techniques and findings offer the possibility to improve the quality control of the manufacturing of periodically poled nonlinear devices, and thus have the potential to improve their performance especially for quantum technology applications.
@@ 2 Experimental setup
The experimental setup for the noise analysis of the frequency converter is shown in Figure 2. The main element of the experiment is the converter: a bulk ppKTP crystal which is polished and coated to also provide a monolithic enhancement cavity for the pump field [29]. We utilise a long flux-grown KTP crystal (Raicol Crystals Ltd.) quasi-phase matched for type-0 DFG (SFG) process 637 nm 1064 nm + 1587 nm (or vice versa) with a poling domain length of , making use of the strong tensor element of the nonlinear susceptibility for the parametric interaction. The conversion bandwidth is 110 GHz (FWHM), corresponding to 0.9 nm (0.15 nm) at 1587 nm (637 nm). The converter is pumped by a commercial 3W CW 1064 nm ND:YAG laser and all measurements presented here were performed at a circulating power of 50 W which yields an internal conversion efficiency . The circulating pump power was inferred from the transmitted pump power and the measured finesse . The monolithic cavity is placed in a oven which is stabilised to less then mK. Further details on the setup and monolithic ppKTP cavity are provided in [29].
@@
The spectra presented in this work were taken with three different spectrometers: two 512 pixel InGaAs spectrometers (OceanInsight NIRQUEST+1.7) with spectral ranges of 900 nm to 1650 nm (resolution 1.82 nm) and 1320 nm to 1490 nm (resolution 0.55 nm). Both have a quantum efficiency of about 0.1% in the standard-gain-mode and about 1% in the high-gain-mode. We also performed a high-resolution measurement over a restricted region of the spectra from 1480 nm to 1620 nm with a resolution of approximately 0.1 nm. This measurement combined a tunable filter (EXFO XTA-50 U) with a transmission of about and a superconducting nanowire single-photon detector (SNSPD) with a quantum efficiency of about 90% and a dark count rate of 100 cps.
@@ 3.1 Measurements of the full spectrum
Figure 3 shows the measurements with the wide-range InGaAs spectrometer from 1140 nm to 1650 nm with a resolution of about 1.82 nm. Below 1140 nm the LPF blocks the remaining pump light. Above 1650 nm the InGaAs line array is no longer sensitive. The four possible combinations of pump and detection polarisation are shown, with each aligned to either the or of the crystal. The blue line shows the measured spectrum for the polarisation configuration corresponding to the type 0 phase-matching condition (both pump and detected mode polarised along the -axis of the crystal). This configuration is also the design configuration of the converter process. Due to the relative strength of the tensor element, this is the configuration where one would anticipate the largest noise floor in the region where parasitic SPDC is the dominant noise source.
The y-axis of Fig. 3 is normalised to the known generated noise spectral density (NSD) of 110 kHz/nm at the target wavelength of 1587 nm [29]. This NSD was measured at 72 W circulating pump power and an internal conversion efficiency of 72%. The crystal temperature was stabilized near with a long-term temperature stability of below 1 mK. All four spectra were corrected for the wavelength-dependent transmission of the optical coating of the monolithic cavity, which was specified by the manufacturer.
In the range from 1140 nm to 1330 nm strong Lorenzian-shaped peaks that we attribute to spontaneous Raman scattering (see Section 3.2) are visible. Beyond 1330 nm, the general trend of decreasing NSD is no longer evident, and an effective pedestal arises that we identify as parasitic SPDC (see Section 3.3). The apparent drop-off beyond 1650 nm is due to the insensitivity of the InGaAs sensor. The succession of narrow-band peaks in the blue graph are due to higher-order counter-propagating SPDC. High resolution measurements of these unique features are presented in Section 3.3. The theoretical background on higher-order counter-propagating SPDC is outlined in Section 4, which was mainly inspired by Fejer et al. [32].
@@ 3.2 Stokes-Raman scattering
In Figure 4 the blue curve () from Figure 3 is plotted over the Raman shift from 500 to 1200 . The Raman shift in terms of wavenumbers is calculated from the spectral wavelength as . The pump wavelength was measured with a wavemeter to be () nm. The resolution of the spectrometer over the plotted range is approximately 13 . The orange curve is a Raman scattering measurement of unpoled KTP extracted from Figure 5 in Ref. [35] from the curve labelled with . The excitation wavelength in Ref. [35] was 632.8 nm.
See Table S1 in the supplemental document for a detailed comparison of the peak positions. Our measurements also qualitatively agree with the Raman measurements in Ref. [36] for bulk KTP and for a proton-implanted KTP waveguide (compare Figure 5 in Ref. [36]) - both z-cut - where an excitation wavelength of 473 nm was used. Further, similar results were presented in Figure 1 of Ref. [37] with an excitation wavelength of 488 nm. As a result, we identify the strong peaks in the range from 1140 nm to 1250 nm (corresponding to a Raman shift of 621 to 1393 ) as Stokes-Raman scattering resonances.
In Figure 5 the noise data was fitted with a sum of Voigt profiles following references [38, 39]. Using this approach, the contributions from the tail of the Raman scattering can be roughly disentangled from the contribution from parasitic SPDC.

### v2

@@ Abstract
Quantum frequency conversion (QFC) will be an indispensable ingredient in future quantum technologies. For example, large-scale fibre-based quantum networks will require QFC to interconnect heterogeneous building blocks like emitters, channels, memories and detectors. The performance of existing QFC devices - typically realised in periodically-poled nonlinear crystals - is often severely limited by parasitic noise that arises when the pump wavelength lies between the inter-converted wavelengths. Here we comprehensively investigate the noise spectrum of a QFC device pumped by a CW 1064 nm laser. The converter was realised as a bulk periodically-poled potassium titanyl phosphate (ppKTP) crystal quasi-phase-matched for conversion between 637 nm and 1587 nm, which was also polished and coated to resonantly enhance the pump field by a factor of 50. While Raman scattering dominates the noise background from 1140 nm to 1330 nm, at larger energy shifts (beyond 60 THz), parasitic spontaneous parametric down-conversion (SPDC) is the strongest noise source. Further, the noise spectrum was contaminated by a regular succession of narrow-band peaks, which we attribute to a heretofore unidentified higher-order counter-propagating SPDC processes - with quasi-phase-matching orders up to 44 evident in our measurements. This work provides a comprehensive overview of the limiting noise sources in QFC devices that use quasi-phase-matched crystals and will prove an invaluable resource in guiding their future development.
@@ 1 Introduction
The efficient and noiseless frequency conversion of quantum light will be an essential building block in the development of future large-scale quantum networks. By enabling the coherent transfer of quantum information between disparate energies, it interfaces the low-loss capabilities of telecommunications infrastructure with the typically visible and near-IR photon energies required by matter-based qubits. One such example are nitrogen-vacancy (NV) centres in diamond [1, 2, 3]. NV centres represent a promising platform due to the long coherence time of the electronic spin associated with this defect centre [4], the possibility to couple this electronic spin to nuclear spin [5, 6, 7] and an optical interface at 637 nm [8]. With NV centres in diamond, quite complex quantum networking tasks have been recently demonstrated [9, 10, 11]. Without the conversion of the 637 nm photons emitted by the NV centre to telecommunication wavelengths, these photons would suffer dramatic transmission losses in optical fibres [12] that limit the size of these networks well below the metropolitan scale. A variety of platforms are similarly limited by this, including other colour centres in diamond like GeV (602 nm), SnV (620 nm), PbV (520/555 nm) [2] or quantum memory at 606 nm [13] and 580 nm [14], trapped ions (369.5 nm) [15] or hexagonal boron nitride (~600 nm) [16].
@@
In these scenarios, quantum frequency conversion (QFC) is often realized as a difference- (sum-) frequency generation process in which a -nonlinear crystal converts a typically short wavelength photon to the telecommunications band (or vice versa) via a comparatively strong pump field at an intermediate wavelength. In this configuration, where the pump wavelength lies between the emitter and target wavelengths, the strong field required to ensure near-unity conversion efficiency also produces considerable unwanted noise at the target wavelength. This parasitic noise is not only present in the conversion from visible to telecommunication wavelengths, but also for the reverse process, since, for an efficient converter, any noise produced will itself be efficiently up- or down-converted. This parasitic noise – whether parametric or non-parametric in origin – has proven to be the limiting factor in the performance of these conversion devices.
@@
Efficient conversion demands a strength of non-linear interaction that is typically inaccessible in a three-wave-mixing process without either waveguide confinement [12], pump enhancement [17] or pulsed pump lasers [18]. Furthermore, frequency converters near-exclusively utilise quasi-phase-matching, as it simultaneously allows for collinear interactions of the fields without spatial walk-off and unlocks access to the largest tensor elements of the nonlinear susceptibility. In theory, this allows for a dramatic reduction in pump power that ideally mitigates noise from non-parametric scattering of the pump photons, notably, Raman scattering and fluorescence. However, errors that arise in the periodic poling can vastly enhance the otherwise strongly suppressed phase-mismatched parasitic spontaneous parametric down-conversion (SPDC) processes. Furthermore, employing waveguide confinement to reduce the needed pump power can lead to a Cerenkov-idler configuration of the SPDC noise process [19] and waveguide inhomogeneities to further undesired phase-matching, both increasing the noise level even more.
@@
This emergent noise floor in the conversion between visible and telecommunication wavelengths was first reported by Albota and Wong [17] who considered conversion in a bulk periodically-poled lithium niobate (ppLN) crystal. Almost two decades on, converters based on ppLN waveguides, which remain the most commonplace platform, are still largely limited by pump-induced noise [20, 21, 22, 23, 24, 12]. The noise spectra of quantum frequency converters based on such devices have been previously investigated in references [25, 24]. In the short-wavelength-pumped scheme of Ref. [24], parasitic SPDC is found to be the dominant source of noise. In the long-wavelength-pumped converter investigated in Ref. [25], anti-Stokes Raman scattering is identified as the main source of noise.
@@
A workaround for this noise problem is two-step conversion. Here the pump wavelength is chosen to be longer than the target wavelength (e.g. 2.1 ) and, due to energy conservation, no unwanted SPDC can be generated at the target wavelength. Additionally, the Raman shift for the anti-Stokes photon is very large, ensuring its noise contribution is negligible. First proposed in [17], two-step conversion was first demonstrated in a single waveguide device [26] and later in two separate waveguides [27]. A drawback of this approach is the necessity for a strong laser source at about 2 alongside the increased complexity of a two-step process. Another promising approach to circumvent the parasitic noise arising in conversion is to utilise critical phase-matching in a bulk crystal [28], eliminating parasitic SPDC noise which would be enhanced by poling errors. Owing to the comparatively smaller nonlinear susceptibility and the effects of double refraction, this approach demands substantially more pump power and the maximum conversion efficiency is limited by the spatial walk-off encountered in critical phase-matching. By utilising external resonant enhancement of the pump, this approach achieved an internal conversion efficiency of 60% for a circulating pump power of 360 W while obtaining a low measured noise of 2.0 kHz/nm. Thus, this approach has about a factor of 20 less noise than a bulk ppKTP converter [29] and two-orders of magnitude less noise then the best ppLN waveguide [12].
@@
In our previous work, we demonstrated an alternative route for low-noise QFC based on a single-step quantum frequency converter based on a monolithic cavity made from bulk periodically-poled potassium titanyl phosphate (ppKTP) [29]. The converter was designed for the conversion of single photons at a wavelength of 637 nm to the target wavelength in the telecommunication L-band (1587 nm) with a strong 1064 nm CW pump laser. Due to the improved poling quality of ppKTP [30, 31], parasitic noise arising from spurious SPDC was reduced by a factor of five compared to the best ppLN waveguide [12], while attaining a high internal conversion efficiency of 72% at about 72 W circulating pump power. The converter, however, still exhibited a pump-induced noise floor of 110 kHz/nm at the target wavelength. Considerable improvements in the noise floor are required for single-step converters based on quasi-phase-matched nonlinear crystals to be a viable technology for frequency conversion between the visible and telecommunication wavelengths.
@@
In this work, we investigate the noise spectrum of the aforementioned bulk ppKTP quantum frequency converter designed to convert photons between 637 nm and 1587 nm via a strong 1064 nm pump laser. Our measurements fully characterise the resulting noise spectrum in the wavelength range from 1140 nm to 1650 nm. By resolving the polarisation and temperature dependence of features within the noise spectrum, we identify different noise contributions and the wavelength regions where they dominate. These results are of immediate relevance for most schemes where the strong pump field lies between the two interconnected wavelengths.
@@
Fig. 1 provides a sketch of our conversion scheme and the primary noise contributions from the different processes. We find that close to this pump wavelength, Raman scattering of pump light is the dominant noise source. As we consider increasingly longer wavelengths approaching the target wavelength, the required phonon energies increase and the contribution of Raman scattering to the noise background is quickly suppressed and spurious spontaneous parametric down-conversion (SPDC) becomes the dominant noise process. This is near-exclusively the case for quasi-phase-matched crystals, where errors in the periodic poling [32] drive spurious SPDC processes that produce a plateau in the noise floor, effectively stable across the transparency of the material. This effect has been studied in detail in [20, 33, 34].
@@
In addition to the anticipated features from Raman scattering and the parasitic SPDC background, we have also identified an exotic nonlinear process - higher-order counter-propagating SPDC - that to our knowledge has never before been experimentally observed. This process, which is only present in periodically-poled devices, can contribute significantly to the parasitic noise background. Notably, both types of counter-propagation, where either the idler or signal photon counter-propagate relative to the pump beam, are observed, with narrow-band peaks of corresponding quasi-phase-matching orders m from 10 up to 44 visible. Owing to poling errors related to the duty cycle, peaks associated with even orders of m are also visible.
@@
In totality, these results provide a comprehensive overview of the noise spectrum of a quantum frequency converter that employs a quasi-phase-matched process and a pump that lies in between the interconnected wavelengths. Our presented measurement techniques and findings offer the possibility to better understand and design periodically-poled nonlinear devices for photonic quantum technologies, e.g. quantum frequency converters for quantum networking.
@@ 2 Experimental setup
The experimental setup for the noise analysis of the frequency converter is shown in Fig. 2. The main element of the experiment is the converter: a bulk ppKTP crystal which is polished and coated to also provide a monolithic enhancement cavity for the pump field [29]. We utilise a long flux-grown KTP crystal (from Raicol Crystals Ltd.) quasi-phase-matched for type-0 DFG (SFG) process 637 nm 1064 nm + 1587 nm (or vice versa) with a poling domain length of , making use of the strong tensor element of the nonlinear susceptibility for the parametric interaction. The conversion bandwidth is 110 GHz (FWHM), corresponding to 0.9 nm (0.15 nm) at 1587 nm (637 nm). The converter is pumped by a commercial 3W CW 1064 nm ND:YAG laser. For practical purposes - ensuring increased stability of the converter - all measurements presented here were performed at a circulating power of 50 W which yields an internal conversion efficiency . The circulating pump power was inferred from the transmitted pump power and the measured finesse . The monolithic cavity is placed in an oven which is stabilised to less than mK. Further details on the setup and monolithic ppKTP cavity are provided in [29].
@@
The spectra presented in this work were taken with three different spectrometers: two 512 pixel InGaAs spectrometers (OceanOptics NIRQUEST+1.7) with spectral ranges of 900 nm to 1650 nm (resolution 1.82 nm) and 1320 nm to 1490 nm (resolution 0.55 nm). Both have a quantum efficiency of about 0.1% in the standard-gain-mode and about 1% in the high-gain-mode. We also performed a high-resolution measurement over a restricted region of the spectra from 1480 nm to 1620 nm with a resolution of approximately 0.1 nm. This measurement combined a tunable filter (EXFO XTA-50 U) with a transmission of about and a superconducting nanowire single-photon detector (SNSPD) with a quantum efficiency of about 90% and a dark count rate of 100 cps (from Quantum Opus LLC).
@@ 3.1 Measurements of the full spectrum
Fig. 3 shows the measurements with the wide-range InGaAs spectrometer from 1140 nm to 1650 nm with a resolution of about 1.82 nm. Below 1140 nm the LPF blocks the remaining pump light. Above 1650 nm the InGaAs line array is no longer sensitive. The four possible combinations of pump and detection polarisation are shown, with each aligned to either the or of the crystal. The blue line shows the measured spectrum for the polarisation configuration corresponding to the type 0 phase-matching condition (both pump and detected mode polarised along the -axis of the crystal). This configuration is also the design configuration of the converter process. Due to the relative strength of the tensor element, this is the configuration where one would anticipate the largest noise floor in the region where parasitic SPDC is the dominant noise source.
The y-axis of Fig. 3 is normalised to the known generated noise spectral density (NSD) of 76 kHz/nm at the target wavelength of 1587 nm at 50 W circulating pump power [29]. The crystal temperature was stabilized near with a long-term temperature stability of below 1 mK. All four spectra were corrected for the wavelength-dependent transmission of the optical coating of the monolithic cavity, which was specified by the manufacturer (Laseroptik GmbH).
In the range from 1140 nm to 1330 nm strong Lorenzian-shaped peaks that we attribute to spontaneous Raman scattering (see Section 3.2) are visible. Beyond 1330 nm, the general trend of decreasing NSD is no longer evident, and an effective pedestal arises that we identify as parasitic SPDC (see Section 3.3). The apparent drop-off beyond 1650 nm is due to the insensitivity of the InGaAs sensor. The succession of narrow-band peaks in the blue graph are due to higher-order counter-propagating SPDC. High resolution measurements of these unique features are presented in Section 3.3. The theoretical background on higher-order counter-propagating SPDC is outlined in Section 4, which was mainly inspired by Fejer et al. [32] and Phillips et al. [34].
@@ 3.2 Stokes-Raman scattering
In Fig. 4 the blue curve () from Fig. 3 is plotted over the Raman shift from 500 to 1200 . The Raman shift in terms of wavenumbers is calculated from the spectral wavelength as . The pump wavelength was measured with a wavemeter to be () nm. The resolution of the spectrometer over the plotted range is approximately 13 . The orange curve is a Raman scattering measurement of unpoled KTP extracted from Fig. 5 in Ref. [35] from the curve labelled with . The excitation wavelength in Ref. [35] was 632.8 nm. See Table S1 in the supplemental document for a detailed comparison of the peak positions. Our measurements also qualitatively agree with the Raman measurements in Ref. [36] for bulk KTP and for a proton-implanted KTP waveguide (compare Fig. 5 in Ref. [36]) - both z-cut - where an excitation wavelength of 473 nm was used. Further, similar results were presented in Fig. 1 of Ref. [37] with an excitation wavelength of 488 nm. As a result, we identify the strong peaks in the range from 1140 nm to 1250 nm (corresponding to a Raman shift of 621 to 1393 ) as Stokes-Raman scattering.
In Fig. 5 the noise data was fitted with a sum of Voigt profiles following references [38, 39]. Using this approach, the contributions from the tail of the Raman scattering can be roughly disentangled from the contribution from parasitic SPDC.